# The Q-AMOEBA (CF) Polarizable Potential

### v2

@@ 3.4 The Alanine Dipeptide
A longstanding question in the field of biomolecular simulation concerns the quantitative and qualitative impacts of NQE 93 in proteins and in protein-ligand systems. Such calculations present inherent challenges due to the high number of conformational states accessible. In this context, we focus on a prototypical system, the alanine dipeptide, a relatively simple yet insightful model that captures key sampling challenges also present in larger proteins. Indeed, it embodies sampling difficulties typical of larger proteins and its well known metastable states characterized by the two dihedral angles and (shown in Figure 3) serves as an ideal model for exploring NQE impacts.
@@
To overcome these, we first resort to the Adaptive Biasing force (ABF) technique 94, 95 using the Colvars library 96 with both the and as collective variables in order to recover the 2 dimensional free energy surface as a function of (,) with NQE using adQTB and without NQE. Details of the simulation, and input files are provided in SI. Our simulations (available in SI) reveal only marginal differences in the free energy surfaces, suggesting that NQE have a minimal effect in this particular system. This highlights the robustness of both the adQTB method and our model in handling complex biological systems. We then investigate the impact of NQE on the hydration free energy of alanine dipeptide using the Q-AMOEBA (CF) model with NQE and its classical counterpart. The results, shown in Table 4, are compared with those from the AMOEBA-03 force field. To compute this, we leverage the recently introduced Lambda-ABF technique within the Tinker-HP framework, coupled to Colvars doi:10.1021/acs.jctc.3c01249. This framework allows for efficient sampling of the alchemical variable while biasing other slow degrees of freedom, such as and . Given that is known to be associated with the largest free energy barrier, we perform 2D ABF simulations with respect to both and , yielding 2D free energy surfaces. The hydration free energy of alanine dipeptide is then recovered by marginalizing with respect to . The hydration free energy computed with the Q-AMOEBA (CF) model and NQE is approximately 4 kcal/mol lower than that obtained using the classical version of the model, indicating that NQE reduces the hydrophilic character of the alanine dipeptide. While electrostatic contributions remain consistent across all models, vdW interactions vary significantly due to the effects of NQE (Figure available in SI). This variation stems from quantum-induced atomic delocalization, particularly in light atoms such as hydrogen, which broadens the range of accessible molecular configurations and enhances entropic contributions. As a result, the strength of vdW interactions changes, while electrostatic interactions—primarily determined by fixed charge distributions and distances—are less influenced. Ultimately, the enhanced sampling enabled by NQE leads to greater entropic contributions to the hydration free energy, causing discrepancies between the quantum and classical models. Interestingly, the AMOEBA-03 results closely matched those of the classical Q-AMOEBA (CF) model, indicating that neglecting NQE, especially in systems where vdW forces play a significant role, can lead to inaccurate free energy estimates.
@@
4 Conclusion
In this work, we introduced Q-AMOEBA (CF), an enhanced version of the Q-AMOEBA model that incorporates a geometry-dependent CF term and explicitly accounts for NQE via the adQTB method. The CF correction addresses key limitations of traditional FF by refining molecular geometries to align bond lengths and angles with experimental data across both gas and liquid phases. Q-AMOEBA (CF) is designed to be used in simulations with explicit inclusion of NQE and provides a step forward towards a more accurate description of molecular properties, particularly in the context of water and biological systems. This enables the prediction of accurate hydration free energies of organic molecules and the study of more complex systems, such as the alanine dipeptide, overall with improved efficiency and precision, showcasing the broad applicability and versatility of Q-AMOEBA (CF). While other models include NQE, their computational expense often limits their use in large systems. The Q-AMOEBA (CF) model offers a balance between accuracy and computational efficiency, marking a milestone in the development of next-generation polarizable force fields for simulating biological and biochemical systems. It is also important to note that the vdW interactions in biological systems may require further attention in order to explicitly use NQE at large scale. This parametrization task will be the subject of further research in link with the biological importance of NQE in biology and will be effectively addressed using machine learning techniques, which hold great potential in training models that adapt the vdW parameters for biological contexts97.

### v3

@@ 0.4 The Alanine Dipeptide
A longstanding question in the field of biomolecular simulation concerns the quantitative and qualitative impacts of NQE 96 in proteins and in protein-ligand systems. Such calculations present inherent challenges due to the high number of conformational states accessible. In this context, we focus on a prototypical system, the alanine dipeptide, a relatively simple yet insightful model that captures key sampling challenges also present in larger proteins. Indeed, it embodies sampling difficulties typical of larger proteins and its well known metastable states characterized by the two dihedral angles and (shown in Figure 3) serves as an ideal model for exploring NQE impacts.
@@
To overcome these, we first resort to the Adaptive Biasing force (ABF) technique 97, 98 using the Colvars library 99 with both the and as collective variables in order to recover the 2 dimensional free energy surface as a function of (,) with NQE using adQTB and without NQE. Details of the simulation, and input files are provided in SI. Our simulations (available in SI) reveal only marginal differences in the free energy surfaces, suggesting that NQE have a minimal effect in this particular system. This highlights the robustness of both the adQTB method and our model in handling complex biological systems. We then investigate the impact of NQE on the hydration free energy of alanine dipeptide using the Q-AMOEBA (CF) model with NQE and its classical counterpart. The results, shown in Table 4, are compared with those from the AMOEBA-03 force field. To compute this, we leverage the recently introduced Lambda-ABF technique within the Tinker-HP framework, coupled to Colvars doi:10.1021/acs.jctc.3c01249. This framework allows for efficient sampling of the alchemical variable while biasing other slow degrees of freedom, such as and . Given that is known to be associated with the largest free energy barrier, we perform 2D ABF simulations with respect to both and , yielding 2D free energy surfaces. The hydration free energy of alanine dipeptide is then recovered by marginalizing with respect to . The hydration free energy computed with the Q-AMOEBA (CF) model and NQE is approximately 4 kcal/mol lower than that obtained using the classical version of the model, indicating that NQE reduces the hydrophilic character of the alanine dipeptide. While electrostatic contributions remain consistent across all models, vdW interactions vary significantly due to the effects of NQE (Figure available in SI). This variation stems from quantum-induced atomic delocalization, particularly in light atoms such as hydrogen, which broadens the range of accessible molecular configurations and enhances entropic contributions. As a result, the strength of vdW interactions changes, while electrostatic interactions—primarily determined by fixed charge distributions and distances—are less influenced. Ultimately, the enhanced sampling enabled by NQE leads to greater entropic contributions to the hydration free energy, causing discrepancies between the quantum and classical models. Interestingly, the AMOEBA-03 results closely matched those of the classical Q-AMOEBA (CF) model, indicating that neglecting NQE, especially in systems where vdW forces play a significant role, can lead to inaccurate free energy estimates.
@@
In this work, we introduced Q-AMOEBA (CF), an enhanced version of the Q-AMOEBA model that incorporates a geometry-dependent CF term and explicitly accounts for NQE via the adQTB method. The CF correction addresses key limitations of traditional FF by refining molecular geometries to align bond lengths and angles with experimental data across both gas and liquid phases. Q-AMOEBA (CF) is designed to be used in simulations with explicit inclusion of NQE and provides a step forward towards a more accurate description of molecular properties, particularly in the context of water and biological systems.
This enables the prediction of accurate hydration free energies of organic molecules and the study of more complex systems, such as the alanine dipeptide, overall with improved efficiency and precision, showcasing the broad applicability and versatility of Q-AMOEBA (CF). While other models include NQE, their computational expense often limits their use in large systems. The Q-AMOEBA (CF) model offers a balance between accuracy and computational efficiency, marking a milestone in the development of next-generation polarizable force fields for simulating biological and biochemical systems. It is also important to note that the vdW interactions in biological systems may require further attention in order to explicitly use NQE at large scale. This parametrization task will be the subject of further research in link with the biological importance of NQE in biology and will be effectively addressed using machine learning techniques, which hold great potential in training models that adapt the vdW parameters for biological contexts100.
Supporting Information
All simulation parameters used in this study, including those for water, ions, and alanine dipeptide, are available in the Supporting Information (SI). The SI also provides the full set of Q-AMOEBA (CF) parameters, along with all input files required for reproducibility and additional figures supporting the results discussed in the main text. Results obtained using TRPMD simulations with Q-AMOEBA (CF), as well as the corresponding water parameters, are provided. In addition, the SI contains the parameters for the Q-AMOEBA (2024) water model (adQTB and TRPMD) together with figures illustrating its performance. Comparative results between Q-AMOEBA (CF), Q-AMOEBA (2024) (TRPMD and adQTB), and the original AMOEBA-03 model are also included. A detailed analysis of ice Ic and heavy water is available. The window parameters used for the hydration free energy calculations, as well as the newly derived Thole parameters for Ca2+ and Mg2+, are provided. Parameters used for all solutes studied are bundled in a tar archive. Finally, for ease of reproducibility, the Colvar input files used for alanine dipeptide simulations are provided, along with the PMF differences obtained with Q-AMOEBA (CF), Q-AMOEBA (2024) (adQTB and PIMD), and AMOEBA-03, as well as the decomposition of the computed solvation free energy for all models.